OpenStudy (aye_my_peeps):

what is 5/16 as a percent?

OpenStudy (ltrout):

Do 5/16 = x/100. Cross mulitply. X = your answer :)

OpenStudy (abdullahm):

To find the percent, you multiply it by 100 and add a % sign net to it.

OpenStudy (abdullahm):

next*

OpenStudy (ltrout):

If you follow mine, the answer should come out to be 31.25%. Do you understand how to do this for future problems?